# Distribution-aware Fairness Learning in Medical Image Segmentation From A Control-Theoretic Perspective

### v1

@@ Abstract
Ensuring fairness in medical image segmentation is critical due to biases in imbalanced clinical data acquisition caused by demographic attributes (e.g., age, sex, race) and clinical factors (e.g., disease severity). To address these challenges, we introduce Distribution-aware Mixture of Experts (dMoE), inspired by optimal control theory. We provide a comprehensive analysis of its underlying mechanisms and clarify dMoE’s role in adapting to heterogeneous distributions in medical image segmentation. Furthermore, we integrate dMoE into multiple network architectures, demonstrating its broad applicability across diverse medical image analysis tasks. By incorporating demographic and clinical factors, dMoE achieves state-of-the-art performance on two 2D benchmark datasets and a 3D in-house dataset. Our results highlight the effectiveness of dMoE in mitigating biases from imbalanced distributions, offering a promising approach to bridging control theory and medical image segmentation within fairness learning paradigms. The source code will be made available.
@@ 3.2 Distribution-aware MoE
Now, we explain our proposed Distribution-aware Mixture of Experts (dMoE), as illustrated in Figure 2(a). We further provide detailed network architecture in Section A.1. Unlike traditional sparse MoE, our proposed dMoE module integrates multiple distribution-wise router networks $G^{attr}$ and a set of $n$ expert modules, consisting of shallow multi-layer perceptron (MLP) neural networks, defined as $E=\{\textit{Expert}_{1},\textit{Expert}_{2},\ldots,\textit{Expert}_{n}\}$. We start by patching intermediate image embeddings from $l$-th layer block output, represented as ${h}_{l}\in\mathbb{R}^{H_{l}W_{l}D_{l}\times Ch_{l}}$, into patches $\tilde{h}_{l}\in\mathbb{R}^{N_{l}\times Ch_{l}}$, where the total number of patches is $N_{l}=H_{l}W_{l}D_{l}$. Here, $H_{l}$, $W_{l}$, $D_{l}$, and $Ch_{l}$ correspond to the height, width, depth, and channel dimensions of the intermediate image embeddings, respectively. For the transformer, which already has a predefined dimension, we omit this process. Given these patched embeddings $\tilde{h}_{l}$ and an attribute flag $attr$, the activated router $G^{attr}$ identifies the top-$k$ experts. The final output is computed as a weighted sum of the outputs of these selected experts:
@@
where $G^{attr}()$ outputs a weight matrix prioritizing each expert’s contribution in a center-specific manner. The resulting weighted output is combined with ${h_{l}}$, representing the shared path, to yield the final unpatched dMoE image embedding $\bar{h_{l}}\in\mathbb{R}^{H_{l}W_{l}D_{l}\times Ch_{l}}$. The router network $G^{attr}$ computes sparse weights $H$ using Gaussian noise, as follows:
@@
$H(x)_{i}=(x^{\top}\cdot W)_{i}+\text{Normal()}\cdot\text{Softplus}((x^{\top}\cdot W^{\text{noise}})_{i}),$ (4)
@@ 3.3 Interpreting dMoE Through Optimal Control
In this section, we begin by providing a brief overview of training deep neural networks from the perspectives of dynamical systems and optimal control. Next, we demonstrate that for a given feedforward architecture as a non-feedback control, MoE can be interpreted as state feedback control. Finally, we show how dMoE relates to the mode-switching control variant of optimal control.
Neural networks create complex transformations by applying a sequence of changes to a hidden state. These transformations can be represented mathematically as follows:
@@
This equation illustrates iterative updates similar to an Euler discretization of a continuous transformation. The dynamics of these transformations are governed by an ODE, parameterized by a neural network, which is a continuous representation of the hidden units’ dynamics:
$dh_{t}=f(h_{t},u_{t})dt.$ (9)
In practical applications, such as training neural networks with a dataset labeled $\{x,y\}$, the goal is to minimize the objective function through an optimization process:
$\begin{split}\arg\min_{\{\theta_{l}\}}&\{\mathbb{E}_{\{x,y\}}L(\hat{y},y)|\\ &h_{l+1}=h_{l}+f(h_{l},\theta_{l}),h_{0}=x,\hat{y}=h_{L}\}.\end{split}$ (10)
This optimization problem can be seen as the discretization of a terminal control problem, which is governed by the neural ODE:
$\begin{split}\arg\min_{\{u^{t}\}}&\{\mathbb{E}_{\{x,y\}}L(\hat{y},y)|\\ &dh_{t}=f(h_{t},u_{t}),h_{0}=x,\hat{y}=h_{L}\}.\end{split}$ (11)
Here, the MoE framework can be considered an approximation of feedback control, also known as closed-loop control, as illustrated in Figure 2(b) (middle). The feedback control ensures optimal operational outcomes by leveraging real-time state information, $h_{t}$ to adjust control inputs dynamically. This is expressed mathematically as:
@@
Given the complexity of optimizing policy function through neural networks, kernel methods (Hofmann et al., 2008) are often employed for parameterization:
$u_{t}(h_{t})\approx\sum_{i}K(h_{t},h_{t}^{i})u^{i}_{t}(h^{i}_{t}),$ (13)
where $K$ is a kernel function, $h_{i}^{t}$ are anchor points, and $u^{i}_{t}(h^{i}_{t})$ are fixed values. This re-parameterization facilitates the shift in optimization focus from the function $u$ itself to the fixed parameters $\theta^{i}_{t}=u^{i}_{t}(h^{i}_{t})$, enhancing the manageability and efficiency of the optimization process:
@@
To facilitate the learning of the kernel function $\phi$ we employ a reparameterization technique where $\langle\phi(h_{t}),\phi(h_{t}^{i})\rangle=\langle h_{t},\phi^{*}\phi(h_{t}^{i})\rangle=(h_{t}^{\top}\cdot W)_{i}$. Here, the transformed parameters $\phi^{*}\phi(h_{t}^{i})$ are represented by the $i$-th column of $W$, with $\phi^{*}$ denoting the transpose of $\phi$. Due to considerations of computational complexity and robustness, we have incorporated Noisy Top-K Gating, corresponding to Eqs. 3, 4, and 5. Specifically, if the function $f$ is linear with respect to its second argument and can be interpreted as the strategy of an expert, where each expert $E_{i}$ acts as a controller with fixed parameters $\theta_{i}$, then the following equation holds:
$\begin{split}f(h_{t},\sum_{i}K(h_{t},h_{t}^{i})\theta^{i}_{t})&=\sum_{i}K(h_{t},h^{i}_{t})f(h_{t},\theta^{i}_{t})\\ &=\sum_{i}G_{i}(h_{t})E_{i}(h_{t}).\end{split}$ (16)
In the field of optimal control, the methodology of choosing or switching between different control strategies based on varying environmental conditions is commonly referred to as mode-switching control or multi-modal control. This approach involves selecting the most suitable control law not only depending on the state $h_{t}$ but also on system parameters $attr$, aiming to optimize overall system performance.
In particular, we consider a special case by introducing a switching logic dependent on the external environmental conditions denoted by $attr$. The dynamics of the system are described by the following equation:
@@
where $\kappa$ is a control strategy mapping that selects the most appropriate control input based on the current state and environment. If the system features several potential control modes, $\kappa$ may comprise multiple sub-strategies, $\kappa_{1},\kappa_{2},\ldots,\kappa_{m}$, each applicable under specific conditions. The logic for switching modes is governed by the switching function $s(attr)$.
To prevent overfitting and to share knowledge learned from images of different distributions, we utilize a shared expert $E_{i}$ across different flags $attr$. For each attribute, a distinct $G$ is trained, implying that for different attr, we adopt different functions $\phi$ in Eq. 15, analogous to using different matrices $W$ in Eq. 4. This results in the following control scheme:
@@
Considering that the neural network’s layerwise structure represents a discrete form of the control, incorporating Patchify leads to Eq. 2. Therefore, dMoE can be interpreted as attribute-wise mode-switching control variant of optimal control for fairness learning, as reflected in the structural resemblance illustrated in Figure 2(a) and (b) (bottom).
@@ 4.3 Baseline Method and Evaluation Metrics
We evaluate our method against baseline approaches for fairness learning. On the 2D Harvard-FairSeg dataset, we compare our approach with reported metrics from adversarial (ADV) training (Madras et al., 2018), distributionally robust optimization (DRO) (Sagawa et al., 2019), fair-error-bound scaling (FEBS) (Tian et al., 2024), and generative model-based augmentation (FairDiff). We implement our method under the same training conditions as the benchmark methods. For further experiments on the 2D HAM10000 dataset and the 3D radiotherapy target dataset, we compare our method with FEBS among SOTA methods. We exclude the generative model-based approach, FairDiff, as it requires additional image generation and evaluation tailored to 2D and 3D datasets, which makes it suboptimal for direct comparison. Moreover, we use the default MoE as an additional baseline, replacing our proposed dMoE modules by default MoE modules without distribution-aware router. We implement all methods under the same training conditions as our method. To evaluate performance, we employ the Dice Similarity Coefficient, Intersection over Union (IoU), as well as equity-scaled segmentation performance (ESSP) metrics for both Dice and IoU—denoted as ES-Dice and ES-IoU—following (Tian et al., 2024):
@@
where $I\in\{\text{Dice},\text{IoU}\}$ and $A$ represents the set of all demographic groups. While ESSP evaluates fairness in performance across diverse demographic subgroups, it has limitations in capturing substantial performance gains for specific subgroups, particularly in the 3D radiotherapy target segmentation task. We complement the limitation of ESSP metrics by providing quantitative analysis with violin plots, which offer a more detailed visualization of equity and performance distribution across diverse subgroups.
@@
ES-Dice Dice ES-IoU IoU Dice IoU Dice IoU Dice IoU
@@
+ MoE 0.834 0.854 0.745 0.762 0.845 0.757 0.842 0.748 0.857 0.765
+ dMoE 0.834 0.862 0.748 0.773 0.844 0.755 0.851 0.761 0.866 0.777
Note. ES-Dice and ES-IoU assess fairness by measuring performance consistency across entire attribute subgroups.
@@
ES-Dice Dice ES-IoU IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice IoU
TransUNet (Chen et al., 2021) 0.831 0.876 0.735 0.824 0.862 0.787 0.868 0.809 0.888 0.846 0.895 0.857 0.875 0.839
+ FEBS (Tian et al., 2024) 0.787 0.858 0.679 0.798 0.831 0.747 0.844 0.774 0.884 0.837 0.871 0.827 0.869 0.830
+ MoE 0.836 0.882 0.742 0.833 0.864 0.794 0.875 0.820 0.889 0.851 0.904 0.869 0.882 0.850
+ dMoE 0.841 0.884 0.749 0.834 0.864 0.791 0.881 0.824 0.890 0.850 0.901 0.866 0.880 0.846
@@
Method All (n=132) T1 (n=1) T2 (n=55) T3 (n=64) T4 (n=12)
ES-Dice Dice ES-IoU IoU Dice IoU Dice IoU Dice IoU Dice IoU
3D ResUNet (Çiçek et al., 2016) 0.583 0.671 0.443 0.521 0.659 0.491 0.635 0.485 0.712 0.564 0.609 0.455
+ FEBS (Tian et al., 2024) 0.549 0.652 0.413 0.496 0.593 0.421 0.587 0.430 0.707 0.554 0.660 0.499
+ MoE 0.584 0.687 0.444 0.540 0.658 0.490 0.611 0.459 0.754 0.614 0.682 0.528
+ dMoE 0.546 0.711 0.424 0.569 0.828 0.707 0.634 0.482 0.765 0.630 0.767 0.626

### v2

@@ Abstract
Ensuring fairness in medical image segmentation is critical due to biases in imbalanced clinical data acquisition caused by demographic attributes (e.g., age, sex, race) and clinical factors (e.g., disease severity). To address these challenges, we introduce Distribution-aware Mixture of Experts (dMoE), inspired by optimal control theory. We provide a comprehensive analysis of its underlying mechanisms and clarify dMoE’s role in adapting to heterogeneous distributions in medical image segmentation. Furthermore, we integrate dMoE into multiple network architectures, demonstrating its broad applicability across diverse medical image analysis tasks. By incorporating demographic and clinical factors, dMoE achieves state-of-the-art performance on two 2D benchmark datasets and a 3D in-house dataset. Our results highlight the effectiveness of dMoE in mitigating biases from imbalanced distributions, offering a promising approach to bridging control theory and medical image segmentation within fairness learning paradigms. The source code is available at https://github.com/tvseg/dMoE.
Machine Learning, ICML
@@ 3.2 Distribution-aware MoE
Now, we explain our proposed Distribution-aware Mixture of Experts (dMoE), as illustrated in Figure 2(a). We further provide detailed network architecture in Section A.1. Unlike traditional sparse MoE, our proposed dMoE module integrates multiple distribution-wise router networks $G^{attr}$ and a set of $n$ expert modules, consisting of shallow multi-layer perceptron (MLP) neural networks, defined as $E=\{\textit{Expert}_{1},\textit{Expert}_{2},\ldots,\textit{Expert}_{n}\}$. We start by flattening intermediate image embeddings from $l$-th layer block output, represented as ${h}_{l}\in\mathbb{R}^{H_{l}W_{l}D_{l}\times Ch_{l}}$, into $\tilde{h}_{l}\in\mathbb{R}^{N_{l}\times Ch_{l}}$, where the total number of embeddings is $N_{l}=H_{l}W_{l}D_{l}$. Here, $H_{l}$, $W_{l}$, $D_{l}$, and $Ch_{l}$ correspond to the height, width, depth, and channel dimensions of the intermediate image embeddings, respectively. For the transformer backbone, which already has a flattened dimension, we omit this process. Given these flattend embeddings $\tilde{h}_{l}$ and an attribute flag $attr$, the activated router $G^{attr}$ identifies the top-$k$ experts. The final output is computed as a weighted sum of the outputs of these selected experts:
@@
where $G^{attr}()$ outputs a weight matrix prioritizing each expert’s contribution in a $attr$-specific manner. The resulting weighted output is combined with $\tilde{h}_{l}$, representing the shared path, to yield the dMoE image embedding $\bar{h_{l}}\in\mathbb{R}^{N_{l}\times Ch_{l}}$. For non-transformer backbones, this representation is reshaped to match spatial dimensions as $\bar{h}_{l}\in\mathbb{R}^{H_{l}W_{l}D_{l}\times Ch_{l}}$, where $N_{l}=H_{l}W_{l}D_{l}$.
In specific, the router network $G^{attr}$ computes sparse weights $H$ using Gaussian noise, as follows:
@@
$H(x)_{i}=(x^{\top}\cdot W)_{i}+\mathcal{N}(0,1)\cdot\text{Softplus}((x^{\top}\cdot W^{\text{noise}})_{i}),$ (4)
@@
Following the dMoE modules, the final dMoE image embedding $\bar{h_{l}}$ is passed to the decoder, which predicts the final output, $\hat{y}$. The network is optimized using the segmentation loss:
@@ 3.3 Interpreting dMoE Through Optimal Control
In this section, we provide a control-theoretic interpretation of deep neural networks by viewing their layer-wise computations as dynamical systems and formulating training as an optimal control problem. This viewpoint offers two key benefits. First, it provides a unified framework for analyzing MoE architectures and yields concrete insights into their adaptive behavior—specifically, how feedback-based control enables more flexible and expressive representations than fixed expert assignments. Second, it facilitates the integration of well-established concepts from control theory into neural network design, including architectural principles, optimization techniques, and regularization strategies. Building on this interpretation, we further show that dMoE naturally corresponds to mode-switching control—a classical paradigm in control theory—and highlight its potential to promote fairness and improve generalization in deep learning systems.
Neural networks create complex transformations by repeatedly updating their internal states (known as hidden features). A common mathematical description of these transformations is given by:
@@
where $h_{l}$ represents the hidden feature vector at layer $l$, and $f(h_{l},\theta_{l})$ represents the transformation applied at layer $l$, parameterized by weights $\theta_{l}$. This iterative updating formula resembles the Euler discretization method used to approximate continuous-time dynamical systems described by ordinary differential equations (ODEs). The continuous-time counterpart of the above discrete transformation can be represented as:
$\frac{dh_{t}}{dt}=f(h_{t},u_{t}),$ (9)
where $h_{t}$ represents the hidden state at continuous time $t$, and $u_{t}$ denotes a control input or parameters guiding the system dynamics at each instant. To clarify further, the discrete layer index $l$ in the neural network corresponds to specific discrete time points sampled from continuous time $t$. Similarly, parameters $\theta_{l}$ at discrete layers are analogous to continuous-time control inputs $u_{t}$, which represent continuously adjustable parameters that influence the evolution of hidden states over time.
In supervised learning applications, the training process of a neural network aims to minimize a loss function $\mathcal{L}$ given labeled data pairs $(x,y)$:
$\begin{split}\arg\min_{\{\theta_{l}\}}&\{\mathbb{E}_{\{x,y\}}\mathcal{L}(\hat{y},y)|\\ &h_{l+1}=h_{l}+f(h_{l},\theta_{l}),h_{0}=x,\hat{y}=h_{L}\}.\end{split}$ (10)
where $\hat{y}$ is the prediction at the final layer $L$, starting from the input $x$. Here, optimizing over parameters $\theta_{l}$ corresponds to choosing the best transformations at each discrete step to minimize the discrepancy between predictions and true labels. From the continuous-time viewpoint, this optimization corresponds to a terminal optimal control problem:
$\begin{split}\arg\min_{\{u^{t}\}}&\{\mathbb{E}_{\{x,y\}}\mathcal{L}(\hat{y},y)|\\ &dh_{t}=f(h_{t},u_{t}),h_{0}=x,\hat{y}=h_{T}\}.\end{split}$ (11)
Here, $\hat{y}$ is the system state at the final time $T$ starting from input $x$ and evolving under control inputs $u_{t}$. Standard feedforward neural networks correspond to Non-feedback control scenarios, meaning the control inputs are determined upfront and do not adapt based on intermediate states. In contrast, the MoE framework introduces adaptive decisions $u_{t}$ that select transformations based on current hidden states $h_{t}$—this is precisely a feedback (closed-loop) control mechanism, as illustrated in Figure 2(b) (middle). The corresponding dynamics are given by:
@@
Given the complexity of optimizing policy function through neural networks, kernel methods (Hofmann et al., 2008) are often utilized to approximate the control function $u_{t}(h_{t})$:
$u_{t}(h_{t})\approx\sum_{i}K(h_{t},h_{t}^{i})u_{t}(h^{i}_{t}),$ (13)
where $K$ is a kernel function, $h_{t}^{i}$ are predefined anchor points, and $u_{t}(h^{i}_{t})$ correspond fixed control parameters at these anchor points. This re-parameterization transforms the optimization process from directly optimizing a potentially complex continuous function $u_{t}$ into optimizing a set of fixed parameters $\theta^{i}_{t}=u^{i}_{t}(h^{i}_{t})$, significantly simplifying the learning task:
@@
To facilitate the learning of the kernel function $\phi$ we introduce a simplified linear reparameterization. The inner product is computed by: $\langle\phi(h_{t}),\phi(h_{t}^{i})\rangle=\langle h_{t},\phi^{*}\phi(h_{t}^{i})\rangle=(h_{t}^{\top}\cdot W)_{i}$, where the transformed anchor features $\phi^{*}\phi(h_{t}^{i})$ form the columns of a learnable parameter matrix $W$, and $\phi^{*}$ denoting the transpose of $\phi$. Furthermore, for computational efficiency and robustness, we have incorporated Noisy Top-K Gating, corresponding to Eqs. (3), (4), and (5). Specifically, if the function $f$ is linear with respect to its second argument and can be interpreted as the strategy of an expert, where each expert $E_{i}$ acts as a controller with fixed parameters $\theta_{i}$, then the following equation holds:
$\begin{split}f(h_{t},\sum_{i}K(h_{t},h_{t}^{i})\theta^{i}_{t})&=\sum_{i}K(h_{t},h^{i}_{t})f(h_{t},\theta^{i}_{t})\\ &=\sum_{i}G_{i}(h_{t})E_{i}(h_{t}),\end{split}$ (16)
where $G_{i}(h_{t})=K(h_{t},h_{t}^{i})$ serve as gating weights dynamically allocating the influence of each expert $E_{i}$. This formulation naturally leads us to mode-switching control, a classical concept in optimal control that aligns with the design philosophy of dMoE. In this framework, the system selects between multiple control strategies based on both the current state $h_{t}$ and external contextual parameters $attr$. Concretely, we model this behavior by introducing a switching logic governed by external conditions $attr$. The system dynamics are defined as:
@@
where $\kappa$ denotes a family of control strategies and $s(attr)$ is a switching function that selects the appropriate mode based on environmental attributes. Each mode $\kappa_{i}$ corresponds to a distinct sub-policy activated under specific conditions. To prevent overfitting and to share knowledge learned from images of different distributions, we utilize a shared expert $E_{i}$ across different flags $attr$. For each attribute, a distinct $G$ is trained, implying that for different attr, we adopt different functions $\phi$ in Eq. (15), analogous to using different matrices $W$ in Eq. (4):
@@
Considering that the neural network’s layerwise structure represents a discrete form of the control leads to Eq. (2). Therefore, dMoE can be interpreted as attribute-wise mode-switching control variant of optimal control for fairness learning, as reflected in the structural resemblance illustrated in Figure 2(a) and (b) (bottom).
@@ 4.3 Baseline Method and Evaluation Metrics
We evaluate our method against baseline approaches for fairness learning. On the 2D Harvard-FairSeg dataset, we compare our approach with reported metrics from adversarial (ADV) training (Madras et al., 2018), distributionally robust optimization (DRO) (Sagawa et al., 2019), fair-error-bound scaling (FEBS) (Tian et al., 2024), and generative model-based augmentation (FairDiff). We implement our method under the same training conditions as the benchmark methods. For further experiments on the 2D HAM10000 dataset and the 3D radiotherapy target dataset, we compare our method with FEBS among SOTA methods. We exclude the generative model-based approach, FairDiff, as it requires additional image generation and evaluation tailored to 2D and 3D datasets, which makes it suboptimal for direct comparison. Moreover, we use the default MoE as an additional baseline, replacing our proposed dMoE modules by default MoE modules without distribution-aware router. We implement all methods under the same training conditions as our method.
To evaluate performance, we employ the Dice Similarity Coefficient, Intersection over Union (IoU), as well as equity-scaled segmentation performance (ESSP) metrics for both Dice and IoU—denoted as ES-Dice and ES-IoU—following (Tian et al., 2024):
@@
where $I\in\{\text{Dice},\text{IoU}\}$ and $A$ represents the set of all demographic groups. ESSP is fundamentally aligned with both the demographic parity and equalized odds principles (Hardt et al., 2016). While ESSP evaluates performance fairness across diverse demographic subgroups, it has limitations in capturing substantial performance gains for specific subgroups, particularly in the 3D radiotherapy target segmentation task. We complement the limitation of ESSP metrics by providing quantitative analysis with violin plots, which offer a more detailed visualization of equity and performance distribution across diverse subgroups. We also report worst-group accuracy metrics to ensure that the model’s lowest performance is properly accounted for.
To support statistical analysis for fairness evaluation, we apply bootstrapped 95% confidence intervals (CIs) when computing both ESSP and all attribute-average metrics, resampling each subgroup and all groups with replacement over 1,000 iterations, respectively.
@@
ES-Dice (CIs) Dice (CIs) ES-IoU (CIs) IoU (CIs) Dice IoU Dice IoU Dice IoU
@@
+ MoE 0.830 (0.809-0.847) 0.854 (0.849-0.860) 0.739 (0.720-0.754) 0.762 (0.755-0.768) 0.845 0.757 0.842 0.748 0.857 0.765
+ dMoE 0.832 (0.810-0.853) 0.862 (0.856-0.867) 0.745 (0.722-0.765) 0.773 (0.766-0.779) 0.844 0.755 0.851 0.761 0.866 0.777
$\dagger$ Metric reported from (Tian et al., 2024). $\ddagger$ ES-metrics are recalculated using Eq. (19), based on metrics reported in the original paper (Li et al., 2024), for a fair comparison.
@@
ES-Dice (CIs) Dice (CIs) ES-IoU (CIs) IoU (CIs) Dice IoU Dice IoU Dice IoU Dice IoU Dice IoU
TransUNet (Chen et al., 2021) 0.792 (0.737-0.841) 0.876 (0.863-0.889) 0.714 (0.664-0.766) 0.824 (0.809-0.838) 0.862 0.787 0.868 0.809 0.888 0.846 0.895 0.857 0.875 0.839
+ FEBS (Tian et al., 2024) 0.757 (0.704-0.807) 0.858 (0.845-0.872) 0.667 (0.613-0.719) 0.798 (0.783-0.812) 0.831 0.747 0.844 0.774 0.884 0.837 0.871 0.827 0.869 0.830
+ MoE 0.796 (0.741-0.844) 0.882 (0.868-0.895) 0.721 (0.671-0.770) 0.833 (0.818-0.846) 0.864 0.794 0.875 0.820 0.889 0.851 0.904 0.869 0.882 0.850
+ dMoE 0.801 (0.745-0.847) 0.884 (0.870-0.896) 0.725 (0.673-0.776) 0.834 (0.820-0.847) 0.864 0.791 0.881 0.824 0.890 0.850 0.901 0.866 0.880 0.846
@@
Method All (n=275) T1 (n=11) T2 (n=129) T3 (n=114) T4 (n=21)
ES-Dice (CIs) Dice (CIs) ES-IoU (CIs) IoU (CIs) Dice IoU Dice IoU Dice IoU Dice IoU
3D ResUNet (Çiçek et al., 2016) 0.487 (0.447-0.529) 0.610 (0.589-0.630) 0.367 (0.336-0.399) 0.462 (0.440-0.482) 0.493 0.341 0.569 0.420 0.659 0.511 0.656 0.506
+ FEBS (Tian et al., 2024) 0.434 (0.406-0.467) 0.586 (0.567-0.604) 0.322 (0.302-0.346) 0.433 (0.414-0.452) 0.442 0.288 0.528 0.374 0.652 0.501 0.685 0.527
+ MoE 0.452 (0.415-0.492) 0.608 (0.586-0.628) 0.342 (0.314-0.372) 0.461 (0.439-0.482) 0.492 0.345 0.542 0.393 0.674 0.532 0.708 0.557
+ dMoE 0.499 (0.469-0.531) 0.650 (0.628-0.671) 0.384 (0.358-0.410) 0.506 (0.483-0.528) 0.718 0.571 0.585 0.435 0.693 0.556 0.778 0.641
Note. The underlined value indicates the worst-group accuracy among distribution attributes for each method.